OpenStudy (anonymous):

Angelo rides his bicycle at an average rate of 14kph. Four hours later, Vance starts after Angelo on a motorcycle and overtakes him in 2 hours, what is Vance's rate?

OpenStudy (anonymous):

Need to setup the problem first: \[4(14) + 2(14) = 2x\] The first 4(14) is what has been done prior to the second person starting. The 2(14) is the time that Angelo is riding. The x is the speed of Angelo. \[84 = 2x\] \[x = 42\] So the answer is 42 kph

OpenStudy (anonymous):

I already came up with that answer but I dont know how to show it in equation. So, thanks for answering. THUMBS UP! :)

OpenStudy (anonymous):

The distance Anglo and Vance ride is the same. Take the distance formula (d=v*t)for each and set them equal to each other. d=distance v=velocity t=time A=Angelo V=Vance so: v(A)*t(A)=v(V)*t(V) and solve for v(V) v(A)*t(a)/t(V)= (14kph*6h)/2h=42kph

OpenStudy (anonymous):

thank you! :)